having trouble solving these equations:
3x-y=3 when x=o
5x-y=7 when x=2
y=-3x-5 when x=1

Answer by ewatrrr(24785)


Hi
3x-y=3 when x=o ⇒ -y = 3 0r y = -3
5x-y=7 when x=2 ⇒ -y = -3 0r y = 3
y=-3x-5 when x=1 ⇒ y = -8